# Asymmetric Errors

### v1

@@ B.15 The linear variance
$\ln L(a)=-{1\over 2}{(a-\hat{a})^{2}\over V+V^{\prime}(a-\hat{a})}$ (83)
@@
$V={\sigma^{+}}{\sigma^{-}}\qquad V^{\prime}={\sigma^{+}}-{\sigma^{-}}$ (84)
@@ B.16 The linear sigma in the log space
Eq. (78) works for any positive, monotonous function $\sigma(a)$ as long as $\sigma(\hat{\alpha}-{\sigma^{-}})={\sigma^{-}}$ and $\sigma(\hat{\alpha}+{\sigma^{+}})={\sigma^{+}}$. As the values of $\sigma$ are supposed to be positive, it seems natural to interpolate these values in the log space. However, simple linear interpolation of $\ln\sigma$ as a function of $a$ is not going to work: for large values of $|a|$ the exponent of a linear function will grow faster than $|a-\hat{\alpha}|$, and the magnitude of $\ln L(a)$ will decrease, creating at least one additional unphysical extremum.
Instead, it can be useful to perform linear interpolation of $\ln\sigma$ as a function of some variable $y=Q(a)$, where $Q(a)$ is a monotonous transform that maps the range of $a$ from [,][$\infty$,$\infty$][ - ∞ , ∞ ] to some compact interval $[y_{\text{min}},y_{\text{max}}]$. Without loss of generality, this interval can be chosen to be $[0,1]$. In this case, $Q(a)$ becomes a cumulative density function of some distribution supported on [,][$\infty$,$\infty$][ - ∞ , ∞ ].
Here, we choose $Q(a)$ to be the cdf of the Fechner distribution whose density is given by Eq. 38 in which we set $m=\hat{a}$, $\sigma_{1}={\sigma^{-}}$, and $\sigma_{2}={\sigma^{+}}$. The equation for $\ln\sigma$ is then
$\ln\sigma(y)=\alpha y+\beta,$ (85)
@@ B.17 The double cubic sigma in the log space
Note that Eq. (78) results in a curve with a continuous second derivative if $\sigma(a)$ is continuous with its first and second derivatives everywhere except the point $a=\hat{a}$. At that point it is sufficient for $\sigma(a)$ just to be continuous, while the continuity of the derivatives is not required (as long as the first and the second derivative are finite on both sides of $\hat{a}$). This allows us to construct a useful log likelihood model as follows:
@@
where $P_{3}^{-}(a)$ and $P_{3}^{+}(a)$ are cubic polynomials satisfying the following conditions: $P_{3}^{-}(\hat{a}-{\sigma^{-}})=\ln{\sigma^{-}}$, $\left.\frac{d}{da}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0$, $\left.\frac{d^{2}}{da^{2}}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0$, $P_{3}^{-}(\hat{a})=P_{3}^{+}(\hat{a})=\ln\sigma_{0}$, $P_{3}^{+}(\hat{a}+{\sigma^{+}})=\ln{\sigma^{+}}$, $\left.\frac{d}{da}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0$, $\left.\frac{d^{2}}{da^{2}}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0$. When the value of the parameter $\sigma_{0}$ is selected, these conditions define the polynomial coefficients uniquely.
@@ B.18 The quintic sigma in the log space
$\ln\sigma(a)=\begin{cases}\ln{\sigma^{-}}&\text{for }a\leq\hat{a}-{\sigma^{-}}\\ P_{5}(a)&\text{for }\hat{a}-{\sigma^{-}}\leq a\leq\hat{a}+{\sigma^{+}}\\ \ln{\sigma^{+}}&\text{for }\hat{a}+{\sigma^{+}}\leq a\end{cases}$ (88)
@@ B.19 The PDG method
In combining results with asymmetric errors the Particle Data Group [PDG] uses ${\sigma^{+}}$ for $a>\hat{a}+{\sigma^{+}}$, ${\sigma^{-}}$ for $a<\hat{a}-{\sigma^{-}}$, and a linearly varying sigma, as given by Equation (80), for the intermediate region. Note that the PDG log likelihood has discontinuous derivatives at $a=\hat{a}-{\sigma^{-}}$ and $a=\hat{a}+{\sigma^{+}}$.
@@ B.21 The skew normal
$\ln L(a)=-{1\over 2}\left({a-\xi\over\omega}\right)^{2}+\ln{\Phi\left(\alpha\left({a-\xi)\over\omega}\right)\right)}$ (90)
@@ Appendix C Implementation in C++/Python
The first package, ”ase”, contains a C++ library implementing probability distributions described in Appendix A and log-likelihood curves described in Appendix B, as well as a number of requisite numerically intensive algorithms (minimization, root finding, special functions, quadratures, convolutions, etc). The URL for this package is https://github.com/igvgit/AsymmetricErrors
The second package, ”asepy”, contains a Python API generated by SWIG [ref:SWIG] for the C++ classes and functions from the first package. It also includes a number of additional high-level utilities and programs for combining errors and results, as well as for plotting distributions and log-likelihoods. This package comes with its own documentation and a sizeable collection of examples. The URL for this package is https://github.com/igvgit/AsymmetricErrorsPy
@@ Appendix D Implementation in R
It comprises a set of functions that use S3 classes to distinguish between different models.
Not all models have been implemented (yet). A list of likelihood models can be found by methods(lnL). There are 4 functions to handle likelihoods:
@@
    takes a list of parameter vectors (which must all be of the same model) and gives the combined result as a parameter vector, though all central values are taken as zero.
@@
    returns the model parameters for a ’flipped’ OPAT result. direction is +1 (the default) when both deviations are positive and -1 when both are negative (the given values of both sp and sm are always positive. Only implemented for the dimidiated model.
Some classes (such as edgeworth and azzalini) are used both for the likelihoods and for the pdfs, but the two are distinct, with no connection.
@@
          legend("topright",col=c("red,green"),lwd=2,
                  legend=c("Dimidiated","Distorted")

### v2

@@ B.15 The linear variance
$\ln L(a)=-{1\over 2}{(a-\hat{a})^{2}\over V+V^{\prime}(a-\hat{a})}.$ (84)
@@
$V={\sigma^{+}}{\sigma^{-}}\qquad V^{\prime}={\sigma^{+}}-{\sigma^{-}},$ (85)
@@ B.16 The linear sigma in the log space
Equation (79) works for any positive, monotonicfunction $\sigma(a)$ as long as $\sigma(\hat{a}-{\sigma^{-}})={\sigma^{-}}$ and $\sigma(\hat{a}+{\sigma^{+}})={\sigma^{+}}$. As the values of $\sigma$ are supposed to be positive, it seems natural to interpolate these values in the log space. However, simple linear interpolation of $\ln\sigma$ as a function of $a$ is not going to work: for large values of $|a|$ the exponent of a linear function will grow faster than $|a-\hat{a}|$, and the magnitude of $\ln L(a)$ will decrease, creating at least one additional unphysical extremum.
Instead, it can be useful to perform linear interpolation of $\ln\sigma$ as a function of some variable $u=Q(a)$, where $Q(a)$ is a monotonictransform that maps the range of $a$ from [,][$\infty$,$\infty$][ - ∞ , ∞ ] to some compact interval $[u_{\text{min}},u_{\text{max}}]$. Without loss of generality, this interval can be chosen to be $[0,1]$. In this case, $Q(a)$ becomes a cumulative density function of some distribution supported on [,][$\infty$,$\infty$][ - ∞ , ∞ ].
Here, we choose $Q(a)$ to be the cdf of the Fechner distribution whose density is given by Equation 36 in which we set $m=\hat{a}$, $\sigma_{1}={\sigma^{-}}$, and $\sigma_{2}={\sigma^{+}}$. The equation for $\ln\sigma$ is then
$\ln\sigma(u)=\alpha u+\beta,$ (86)
@@ B.17 The double cubic sigma in the log space
Note that Equation (79) results in a curve with a continuous second derivative if $\sigma(a)$ is continuous with its first and second derivatives everywhere except the point $a=\hat{a}$. At that point it is sufficient for $\sigma(a)$ just to be continuous, while the continuity of the derivatives is not required (as long as the first and the second derivative are finite on both sides of $\hat{a}$). This allows us to construct a useful log likelihood model as follows:
@@
where $P_{3}^{-}(a)$ and $P_{3}^{+}(a)$ are cubic polynomials satisfying the following conditions:
$P_{3}^{-}(\hat{a}-{\sigma^{-}})=\ln{\sigma^{-}},$
$\left.\frac{d}{da}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{3}^{-}(a)\right|_{a=\hat{a}-{\sigma^{-}}}=0,$
$P_{3}^{-}(\hat{a})=P_{3}^{+}(\hat{a})=\ln\sigma_{0},$ (89)
$P_{3}^{+}(\hat{a}+{\sigma^{+}})=\ln{\sigma^{+}}.$
$\left.\frac{d}{da}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0,$
$\left.\frac{d^{2}}{da^{2}}P_{3}^{+}(a)\right|_{a=\hat{a}+{\sigma^{+}}}=0.$
@@ B.18 The quintic sigma in the log space
$\ln\sigma(a)=\begin{cases}\ln{\sigma^{-}}&\text{for }a\leq\hat{a}-{\sigma^{-}}\\ P_{5}(a)&\text{for }\hat{a}-{\sigma^{-}}\leq a\leq\hat{a}+{\sigma^{+}}\\ \ln{\sigma^{+}}&\text{for }\hat{a}+{\sigma^{+}}\leq a\end{cases}.$ (90)
@@ B.19 The PDG method
In combining results with asymmetric errors the Particle Data Group [28] uses ${\sigma^{+}}$ for $a>\hat{a}+{\sigma^{+}}$, ${\sigma^{-}}$ for $a<\hat{a}-{\sigma^{-}}$, and a linearly varying sigma, as given by Equation (81), for the intermediate region. Note that the PDG log likelihood has discontinuous derivatives at $a=\hat{a}-{\sigma^{-}}$ and $a=\hat{a}+{\sigma^{+}}$.
@@ B.21 The skew normal
$\ln L(a)=-{1\over 2}\left({a-\xi\over\omega}\right)^{2}+\ln{\Phi\left(\alpha\left[{a-\xi)\over\omega}\right]\right)}$ (92)
@@ Appendix C Implementation in C++/Python
The first package, “ase”, contains a C++ library implementing probability distributions described in Appendix A and log-likelihood curves described in Appendix B, as well as a number of requisite numerically intensive algorithms (minimization, root finding, special functions, quadratures, convolutions, etc). The URL for this package is https://github.com/igvgit/AsymmetricErrors
The second package, “asepy”, contains a Python API generated by SWIG [29] for the C++ classes and functions from the first package. It also includes a number of additional high-level utilities and programs for combining errors and results, as well as for plotting distributions and log-likelihoods. This package comes with its own documentation and a sizeable collection of examples. The URL for this package is https://github.com/igvgit/AsymmetricErrorsPy
@@ Appendix D Implementation in R
The package comprises a set of functions that use S3 classes to distinguish between different models. Not all models have been implemented (yet). A list of likelihood models can be found by methods(lnL).
There are 4 functions to handle likelihoods:
@@
    takes a list of parameter vectors, which must all be of the same model, and gives the combined result as a parameter vector. (The central value of the result, which is set to zero, is meaningless.)
These use named vectors as parameters and return values. The indexing of parameters within these vectors is not standardised: access should always be by name.
@@
    returns the model parameters for a ’flipped’ OPAT result. direction is $+$1 (the default) when both deviations are positive and --1 when both are negative (the given values of both sp and sm are always positive. Only implemented for the dimidiated model.
@@
Some classes (such as edgeworth and azzalini) are used both for the likelihoods and for the pdfs, but the two are distinct, with no connection. The function Title(p), which returns the name of the model of the parameters in printable form, with spaces and upper-case letters, may be useful.
@@
          legend("topright",col=c("red","green"),lwd=2,
                 legend=c("Dimidiated","Distorted"))
@@
  • Likelihood and Pdf Errors
    1. 1.
      The combined results of Example 5.3 were produced by the code
      poissonerrors <- function(x){
          L0 <- x*log(x)-x
          g <- function(y) { return (-y+x*log(y)- L0 +.5)}
          hi <- uniroot(g,c(x,x+2*sqrt(x)),tol=1.e-8)
          lo <- uniroot(g,c(x,max(0.0001,x-2*sqrt(x))),tol=1.e-8)
          return(c(sp=hi\$root-x,sm=x-lo\$root))
      phi <- function(x){ # seek Delta ln L = -1/2 upwards
         xx <- x[1]
         L0 <- x[2]
         g <- function(y){  return(-y+xx*log(y)-L0+.5)}
         return(uniroot(g,c(xx,xx+2*sqrt(xx)),tol=1.e-8)\$root-xx)
      plo <- function(x){ # seek Delta ln L = -1/2 downwards
         xx <- x[1]
         L0 <- x[2]
         g <- function(y){  return(-y+xx*log(y)-L0+.5)}
         return(xx-uniroot(g,c(xx,max(0.0001,xx-2*sqrt(xx))),tol=1.e-8)\$root)
      model1 <- "distorted"
      model2 <- "linearvariance"
      t <- 1 # recording time of 1 second
      s <- as.numeric(readline("Error on time in sec - suggest 0.1 "))
      sp <- 1/(t-s)-1/t
      sm <- 1/t-1/(t+s)
      rt <- 1/t
      print(paste("result for time factor ",rt," + ",sp, " - ",sm))
      A <- 100 # acceptance factor
      n <- as.numeric(readline("Expected number of events - suggest 50 "))
      pp <- poissonerrors(n)
      print(paste(" Number with poisson errors is ",n,"+",pp[’sp’],"-",pp[’sm’]))
      r <- n*A*rt # the result
      sp1 <- as.numeric(pp[’sp’]*A*rt)
      sm1 <- as.numeric(pp[’sm’]*A*rt)
      sp2 <- n*A*sp
      sm2 <- n*A*sm
      print(paste("Quoted result ", r, "+",sp1,"-",sm1," (stat) +",sp2,"-",sm2," (syst)"))
      p1p <- getPdfpars(c(M=r,sp=sp1,sm=sm1),model1)
      p2p <- getPdfpars(c(M=r,sp=sp2,sm=sm2),model1)
      p3p <- combinePdferrors(list(p1p,p2p))
      print("Combine (wrongly) as two pdfs"); print(p3p)
      p1L <- getlnLpars(c(val=r,sp=sp1,sm=sm1),model2)
      p2L <- getlnLpars(c(val=r,sp=sp2,sm=sm2),model2)
      p3L <- combinelnLerrors(list(p1L,p2L))
      print("combine(wrongly) as two likelihoods"); print(p3L)